Fermi National Laboratory
 

Accelerator Update for February 28, 2007 - March 2, 2007

 

Wednesday February 28

The day shift began with Operations monitoring the Recycler stash (430.8E10), with the Antiproton Source (Pbar) stacking (21.8E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam in its horn off mode, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the Tevatron (TeV) into shot setup around 8:15 AM.

At 9:31 AM, the TeV quenched during a dry squeeze at sector B0.

Operations prepared the CDF and D0 collision halls for an access at 10:26 AM.

TeV experts reported at 10:56 AM that a control card for a TeV power supply (QDD1) was the cause of the quench. They replaced the card.

At 11:26 AM, a Booster RF station (BRF6) tripped off. It reset.

D0 came out of access at 11:35 AM.

Cryo system experts reported at 12:25 PM that they had completed adjusting the exhaust valves for the TeV sector D1 wet engine.

At 12:30 PM, Operations reported that the MI beam permit tripped off due to beam on the MI-8 collimators for the first of three times in six minutes. A Booster expert reported that he was tuning the Booster beam and that was causing the problems.

At 12:33 PM, NuMI tripped off due to a quadrupole power supply trip (QD108).

CDF came out of access at 12:49 PM.

Operations put the TeV through a dry squeeze at 1:41 PM.

TeV experts began a sextupole study at 2:10 PM.

Operations reported at 4:09 PM that the Recycler had stashed antiprotons three times during the day shift.

Operations reported that NuMI horn experts had the horn up and running by 5:25 PM.

At 6:45 PM, the NuMI beam permit tripped off due to a horn trip. It reset.

Operations put the TeV through a dry squeeze at 7:20 PM.

Operations put the TeV into shot setup at 7:53 PM

At 7:44 PM, the NuMI beam permit tripped off due to a horn trip. It reset.

At 10:10 PM, the TeV suffered a sector D4 quench just before the store was to be established. Operations and TeV experts investigated, but they couldn’t find a cause for the quench.

At 11:52 PM, the NuMI beam permit tripped off due to a horn trip. It reset.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Thursday March 1

The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (12.2E10), with Pbar stacking (104.0E10), and with NuMI taking beam.

Operations ramped the TeV at 1:43 AM.

Operations put the TeV into a dry squeeze at 1:59 AM and an expert began a study.

At 2:50 AM, the NuMI horn started tripping off. Operations contacted an expert.

At 3:04 AM, Operations turned off the NuMI beam while experts decided what to do.

A power supply expert told operators to adjust the current for the horn’s power supply. NuMI resumed taking beam at 3:56 AM.

At 5:01 AM, the NuMI horn started tripping again. Operations took beam away from NuMI.

TeV studies ended at 5:05 AM.

Operations put NuMI into its horn off running mode at 5:20 AM.

Operations put the TeV into shot setup at 7:17 AM.

Operations reported at 7:54 AM that the Recycler had stashed twice during the midnight shift.

Operations reported at 8:49 AM that Pbar experts had switched to a different spot on the antiproton target.

Operations established store 5256 at 10:11 AM with an initial luminosity of 133.5E30.

At 10:26 AM, LCW technicians received approval to access Meson M01 to search for a LCW leak. They reported that they found and repaired a leak on an MWest magnet (MW1W-3).

Operations reported at 12:06 PM that the NuMI horn was back on and taking beam.

At 1:15 PM, the MI-8 collimators pulled the MI abort for the first of seven times in three hours. Everything reset each time.

At 3:05 PM, NuMI tripped off due to a bad beam position on the target. It reset.

At 3:25 PM, the MI beam permit tripped off due to bad beam positions.

Operations reported at 4:22 PM that the Recycler had stashed three times during the day shift.

At 5:56 PM, the MI beam permit tripped off for the first of three times in three hours due to the MI-8 collimators. It reset each time.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Friday March 2

The midnight shift began with Operations monitoring store 5256 (36.94E30) and the stash (237.4E10), with Pbar stacking (31.4E10), and with NuMI taking beam.

At 12:23 AM, the MI beam permit tripped off due to the MI-8 collimators. It reset.

Operations terminated store 5256 at 2:03 AM.

Operations put the TeV into shot setup at 2:20 AM.

At 4:27 AM, the TeV quenched at sector B0 just before the next store was to be established. Operators and experts investigated.

TeV experts began a study period at 6:08 AM.

A Switchyard expert began a one-hour study at 6:59 AM.

Operations reported at 7:54 AM that a door in the TeV sector D had gone unsecure and dropped the area’s interlock, which caused the TeV to quench again. Operators contacted the Interlock techs.

Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to find out why the TeV tunes are ending up in the wrong places. The Run Coordinator will allow Linac experts to work on LRF4 later today and to let Booster experts conduct a study. Pbar and Recycler need to stack and stash so we can establish a store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/05/2007    email Fermilab